# Higher-Order Regularization Learning on Hypergraphs

### v1

@@ 5.1 Active learning
Optimization problems of the form $\arg\min_{v}J(v)+\Psi(v,y)$, where $J$ is a regularizer and $\Psi$ enforces label fidelity, admit a Bayesian interpretation. Specifically, with a prior $\mu_{0}(v)$ proportional to $e^{-J(v)}$, a likelihood $\mu_{1}(y|v)$ proportional to $e^{-\Psi(v,y)}$, we obtain a posterior $\mu_{2}(v|y)$ that is proportional to $e^{-J(v)-\Psi(v,y)}$ implying that the maximum à posteriori estimator of $\mu_{2}$ is the minimizer of $J(v)+\Psi(v,y)$. This formulation enables uncertainty quantification and active learning, see [97, 45, 58].
@@
In graph-based approaches, the regularizer is often chosen as $J(v)=\langle v,L^{s}v\rangle_{n}$ for some $s>0$, where $L$ is the graph Laplacian [96, 22, 58, 82]. This choice induces a Gaussian prior over functions, leveraging the fact that $L$ is a symmetric and positive semi-definite matrix [82]. By Proposition 3.3, an analogous construction is possible on hypergraphs using the operator $\mathcal{L}^{(q)}_{\mathrm{Dis}}$, allowing us to define Gaussian priors in the hypergraph setting as well. This introduces higher-order structure into the prior, effectively encoding regularity up to the $p_{q}$-th derivative [85].
We evaluate this approach within an active learning setting, employing uncertainty sampling as the acquisition function [70]. Experiments are conducted on the MNIST [52] and FashionMNIST [89] datasets. Since both datasets can be embedded in metric spaces, we approximate HOHL by (3) and, following standard practice to speed-up computation on large datasets  [11], we construct $k$-nearest neighbor graphs instead, replacing the scale sequence $\varepsilon^{(\ell)}$ in Eq. (3) with neighborhood sizes $k^{(1)}\geq\dots\geq k^{(q)}$. Edge weights are defined by $w_{k^{(\ell)},ij}=\exp\left(-\frac{4\|x_{i}-x_{j}\|^{2}}{d_{k^{(\ell)}}(x_{i})^{2}}\right),$ where $d_{k^{(\ell)}}(x_{i})$ is the distance from $x_{i}$ to its $k^{(\ell)}$-th nearest neighbor. We choose the norm $\|\cdot\|$ to be the cosine/angular distance.
@@ 5.2 HOHL for semi-supervised learning in non-geometric setting
We consider the Zoo [21], Mushroom [21], Cora [56] and Citeseer [69] datasets. The hyperedges are created following the procedure detailed in Section 3.4.3. To ease notation, in this section, we will write $\mathcal{L}^{(q)}$ instead of $\mathcal{L}^{(q)}_{\mathrm{Dis}}$.
@@
We compare against Laplace Learning using the clique expansion—chosen over other hypergraph-to-graph reductions for its preservation of the vertex set, see [90]—as well as three non-deep hypergraph methods implemented in [30]: transductive learning from [93], hyperedge-weighted transduction from [29], and dynamic hypergraph learning from [92]. We report mean accuracies and standard deviation in percentages over 100 trials at different labelling rates in Tables 3, 4, 5 and 6. We summarize the terminology used in our experiments in Table 2. Similar experiments have been performed to test HOHL in the geometric setting in [85].
@@ 6 Conclusion
Future work includes analyzing HOHL through the lens of reproducing kernel Hilbert space (RKHS) theory, following approaches such as [91], to derive expected error bounds in the semi-supervised setting as a function of the length-scales. Additionally, adaptive skeleton segmentation and parameter selection strategies—e.g., cross-validation, meta-learning, or Bayesian optimization—could further improve robustness. Finally, integrating HOHL into end-to-end differentiable models may enable closer connections to neural architectures, while extending it to dynamic or multilayer hypergraphs opens avenues for application to temporal and multiplex data.

### v2

@@ 5.1 Active learning
Optimization problems of the form $\arg\min_{v}J(v)+\Psi(v,y)$, where $J$ is a regularizer and $\Psi$ enforces label fidelity, admit a Bayesian interpretation. Specifically, with a prior $\mu_{0}(v)$ proportional to $e^{-J(v)}$, a likelihood $\mu_{1}(y|v)$ proportional to $e^{-\Psi(v,y)}$, we obtain a posterior $\mu_{2}(v|y)$ that is proportional to $e^{-J(v)-\Psi(v,y)}$ implying that the maximum à posteriori estimator of $\mu_{2}$ is the minimizer of $J(v)+\Psi(v,y)$. This formulation enables uncertainty quantification and active learning, see [LapRefActiveLearning, ji12, Miller].
@@
In graph-based approaches, the regularizer is often chosen as $J(v)=\langle v,L^{s}v\rangle_{n}$ for some $s>0$, where $L$ is the graph Laplacian [LapRef, Stuart, Miller, TutSpec]. This choice induces a Gaussian prior over functions, leveraging the fact that $L$ is a symmetric and positive semi-definite matrix [TutSpec]. By Proposition 3.4, an analogous construction is possible on hypergraphs using the operator $\mathcal{L}^{(q)}_{\mathrm{Dis}}$, allowing us to define Gaussian priors in the hypergraph setting as well. This introduces higher-order structure into the prior, effectively encoding regularity up to the $p_{q}$-th derivative [weihs2025Hypergraphs].
We evaluate this approach within an active learning setting, employing uncertainty sampling as the acquisition function [settles2012active]. Experiments are conducted on the MNIST [LeCun1998] and FashionMNIST [xiao2017fashionmnist] datasets. Since both datasets can be embedded in metric spaces, we approximate HOHL by (3) and, following standard practice to speed-up computation on large datasets  [poissonLearning], we construct $k$-nearest neighbor graphs instead, replacing the scale sequence $\varepsilon^{(\ell)}$ in Eq. (3) with neighborhood sizes $k^{(1)}\geq\dots\geq k^{(q)}$. Edge weights are defined by $w_{k^{(\ell)},ij}=\exp\left(-\frac{4\|x_{i}-x_{j}\|^{2}}{d_{k^{(\ell)}}(x_{i})^{2}}\right),$ where $d_{k^{(\ell)}}(x_{i})$ is the distance from $x_{i}$ to its $k^{(\ell)}$-th nearest neighbor. We choose the norm $\|\cdot\|$ to be the cosine/angular distance.
@@ 5.2 HOHL for semi-supervised learning in non-geometric setting
We consider the Zoo [Dua:2017], Mushroom [Dua:2017], Cora [McCallum2000] and Citeseer [Sen2008] datasets. The hyperedges are created following the procedure detailed in Section 3.4.3. To ease notation, in this section, we will write $\mathcal{L}^{(q)}$ instead of $\mathcal{L}^{(q)}_{\mathrm{Dis}}$.
@@
We compare against Laplace Learning using the clique expansion—chosen over other hypergraph-to-graph reductions for its preservation of the vertex set, see [clique]—as well as three non-deep hypergraph methods implemented in [hgLearningPractice]: transductive learning from [scholkopfHyper2006], hyperedge-weighted transduction from [hgWeighting], and dynamic hypergraph learning from [dynamicHG]. We report mean accuracies and standard deviation in percentages over 100 trials at different labelling rates in Tables 3, 4, 5 and 6. We summarize the terminology used in our experiments in Table 2. Similar experiments have been performed to test HOHL in the geometric setting in [weihs2025Hypergraphs].
@@ 6 Conclusion
Future work includes analyzing HOHL through the lens of reproducing kernel Hilbert space (RKHS) theory, following approaches such as [zhang2005spectral], to derive expected error bounds in the semi-supervised setting as a function of the length-scales. Additionally, adaptive skeleton segmentation and parameter selection strategies—e.g., cross-validation, meta-learning, or Bayesian optimization—could further improve robustness. Finally, integrating HOHL into end-to-end differentiable models may enable closer connections to neural architectures, while extending it to dynamic or multilayer hypergraphs opens avenues for application to temporal and multiplex data.
Acknowledgments
AW and AB were supported in part by NSF grant DMS-2152717. MT acknowledges the support of the EPSRC Mathematical and Foundations of Artificial Intelligence Probabilistic AI Hub (grant agreement EP/Y007174/1), the Leverhulme Trust through the Project Award “Robust Learning: Uncertainty Quantification, Sensitivity and Stability” (grant agreement RPG-2024-051) and the NHSBT award 177PATH25 “Harnessing Computational Genomics to Optimise Blood Transfusion Safety and Efficacy”.